OpenStudy (surana):

Given b=2 and h = 8, what is the equation of the graph if the parent function is y = /- positive x?

OpenStudy (surana):

The /- is meant to represent square root, or basically it should look like this\[\sqrt{x}\]

OpenStudy (surana):

And the possible answers are a: y = \[\sqrt{2x+16}\] B: y = \[\sqrt{2x-16}\] C: y = \[\sqrt{2x}+16 \] D: y = \[\sqrt{2x} - 16 \]

zepdrix (zepdrix):

I don't know what b represents in this case. We usually use the letter a for amplitude but I guess that's not what the b is for in this case. What does your standard form look like?\[\Large\rm y=\sqrt{b(x-h)}+k\]\[\Large\rm y=\sqrt{bx-h}+k\]One of those maybe?

zepdrix (zepdrix):

Yah I guess it's that first one that I posted, that's the only one which leads to a possible answer. Remember that h represents a `horizontal shift`, so that one needs to get stuck under the root with the x.

OpenStudy (surana):

B is 2.

OpenStudy (surana):

Unless you meant the answer of B.

zepdrix (zepdrix):

\[\Large\rm y=\sqrt{b(x-h)}\]\[\Large\rm y=\sqrt{2(x-8)}\]B is 2?

OpenStudy (surana):

B = 2 in terms of the variable.

zepdrix (zepdrix):

? 0_o

OpenStudy (surana):

Wait a minute...So technically it would be 2x+16, so that would mean A?

OpenStudy (anonymous):

i have seen this before this is not math this is some system or math teacher made up math

zepdrix (zepdrix):

oh b=2, not B=2. Yah that's why I was getting confused by what you were saying :c heh

OpenStudy (surana):

Whoops. I should have been more specific.

zepdrix (zepdrix):

Yah it seems like this relates to some specific system that a teacher didn't tell the rest of us about :3 lol

zepdrix (zepdrix):

A lil tricky to figure out

OpenStudy (anonymous):

what is \(b\)? what is \(h\)? who knows there is nothing here remotely in the realm of mathematics there is no standard notation that says \(h\) is something and \(b\) is something, it it utter nonsense refuse to do it

OpenStudy (anonymous):

even the concept of "parent function" is silly

OpenStudy (surana):

Aren't those variables? And b is equal to 2 and h is equal to 8.

OpenStudy (anonymous):

yeah but what do they represent?

zepdrix (zepdrix):

Nahhh don't listen to that grump ^ We can sort of read into what the teacher probably meant for this problem lol. You want to `subtract` your h value. So we're looking for something that looks like this probably:\[\Large\rm y=\sqrt{b(x-h)}\]

OpenStudy (anonymous):

ooh i wouldn't bet on it !!

zepdrix (zepdrix):

If we plug in our 2 and our 8,\[\Large\rm y=\sqrt{2(x-8)}\]Does that look like any of your options? Can we do something to match it up? :)

OpenStudy (anonymous):

could be \(b(f(x)+h\) or \(f(bx+h)\) o r\(f(b(x+h))\) or \(f(bx-h)\)

OpenStudy (surana):

I'm pretty sure that we could multiply 2 into x and -8.

OpenStudy (surana):

Which would mean 2x - 16

zepdrix (zepdrix):

ya that's probably what teach was looking for! :) So option B, ya?

OpenStudy (surana):

B? Yeah.

OpenStudy (surana):

Regarding that other guy who was here, is that normal of him?

zepdrix (zepdrix):

Yes. He gets grumpy over teachers mistakes and unconventional teaching methods :) lol He's a silly guy.

OpenStudy (surana):

Okay, I guess the "if you don't have anything constructive to say then leave" response will be unnecessary.

zepdrix (zepdrix):

Ehh he's just gotta get it out of his system sometimes XD I dunno, its weird haha

OpenStudy (surana):

Alright. I'll let it alone then. Thanks.